Write the equation of the line in slop-intercept form of the line that is perpendicular to y= -5/3x - 2 and passes through (-3,-3).
Answer by venugopalramana(3286)

Write the equation of the line in slop-intercept form of the line that is perpendicular to y= -5/3x - 2 and passes through (-3,-3).
PERPENDICULAR IS GIVEN BY
Y=3X/5+K
IT PASSES THROUGH (-3,-3)...SO..
-3=3*-3/5+K
K=9/5 -3=--6/5
Y=3X/5-6/5